The area of a square is (16x^2 + 40x +25cm^2) what is x if the perimeter of the square is 52cm?
Answer by ankor@dixie-net.com(22740)

The area of a square is (16x^2 + 40x +25cm^2) what is x if the perimeter of the square is 52cm?
:
P = 52
therefore one side of the square (s)
s = 52/4
s = 13 cm
:
A = s^2
therefore
A = 13^2
A = 169 sq/cm
:
So we have:
16x^2 + 40x + 25 = 169
16x^2 + 40x + 25 - 169 = 0
16x^2 + 40x - 144 = 0
Simplify, divide by 8
2x^2 + 5x - 18 = 0
Factors to
(2x+9)(x-2) = 0
positive solution
x = 2
;
;
See if that checks out, replace x with 2 in the given expression
16(2^2) + 40(2) + 25
64 + 80 + 25 = 169